Please help w/ the following. Thank You.

If 10,000 is invested at 9% compounded monthly for 6 years the compounded amount is given by A=10,000(1.0075)^72. Given that log 1.0075=0.00325, find log A. (Note that 10,000=10^4).

Answer by stanbon(57262)

If 10,000 is invested at 9% compounded monthly for 6 years the compounded amount is given by A=10,000(1.0075)^72. Given that log 1.0075=0.00325, find log A.
------------
log A = log 10000 + 72(log 1.0075)
log A = 4 +72(0.00325)
log A = 4 + 0.234
log A = 4.234
Cheers,
Stan H.